OpenStudy (anonymous):

9-y^2/3-y multiplied by 12y/9+3y????

OpenStudy (anonymous):

\[\frac{9-y^{2}}{3-y}*\frac{12y}{9+3y}\] Is this what you mean?

OpenStudy (anonymous):

exacty!

OpenStudy (anonymous):

Just do them all. 12y*9 minus 12y*y^2 divided by 3*9+3*3y-9*y-y*3y and simplify if you can.

OpenStudy (anonymous):

\[ \frac{9\times y^{2} }{ 3\times9+3\times3y-9\times y-y \times 3y} \] is that what you meant?

OpenStudy (anonymous):

Not sure, I haven't solved it. The first two operations I left indicated go on top, all the other stuff goes on the bottom...good luck.

OpenStudy (anonymous):

haha, thanks it was 4y

OpenStudy (anonymous):

\[\frac{108y-12y^3}{3×9+3×3y−9×y−y×3y}\] after you solve the bottom part, try to factor out some stuff to simplify.

OpenStudy (anonymous):

the answer to the whole problem was 4y, i already solved it thank you for your help.

OpenStudy (anonymous):

yw.